OpenStudy (anonymous):

(6+sqrt-81)(7+sqrt-9) ... gahh once i think i understand something?!

OpenStudy (anonymous):

in an a+bi

OpenStudy (anonymous):

(6+9i)(7+3i)= 42 +63i+18i-27=15+81i

OpenStudy (anonymous):

thank you!